Quantitative Problem: Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis. Both projects' after-tax cash flows are shown on the time line below. Depreciation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firm's average project. Bellinger's WACC is 10%.

YEARS 0 1 2 3 4

Project A -1,000 650 355 290 340

Project B -1,000 250 290 440 790

What is Project A's payback? Round your answer to four decimal places. Do not round your intermediate calculations.

_____years

What is Project A's discounted payback? Round your answer to four decimal places. Do not round your intermediate calculations.

_____years

What is Project B's payback? Round your answer to four decimal places. Do not round your intermediate calculations.

____years

What is Project B's discounted payback? Round your answer to four decimal places. Do not round your intermediate calculations.

______years
Answer #1
Year Cash flows Cumulative cash flows PVF PV of cash flows Cumulative Present value
0 -1000 -1000 1 -1000 -1000
1 650 -350 0.909090909 590.9090909 -409.0909091
2 355 5 0.826446281 293.3884298 -115.7024793
3 290 295 0.751314801 217.8812923 102.1788129
4 340 635 0.683013455 232.2245748 334.4033877
Payback period = 1 + 350/355 = 1.99 years
Discounted payback period =2+115.7/217.88 = 2.53 years
Project B
Year Cash flows Cumulative cash flows PVF PV of cash flows Cumulative Present value
0 -1000 -1000 1 -1000 -1000
1 250 -750 0.909090909 227.2727273 -772.7272727
2 290 -460 0.826446281 239.6694215 -533.0578512
3 440 -20 0.751314801 330.5785124 -202.4793388
4 790 770 0.683013455 539.5806297 337.1012909
Payback period = 3 + 20/790 = 3.03 years
Discounted payback period =3+202.48/539.58 = 3.38 years